# Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation

### v1

@@ 2.2 Fixed-effect meta-analysis (Stouffer’s method)
where $\sigma_{\text{MA}}^{2}=1/(1/\sigma_{1}^{2}+1/\sigma_{2}^{2})$ and
@@
The first equation in (8) represents Stouffer’s p-value combination method (after transforming p-values to $z$-values) using inverse standard errors as weights [19], whereas the second equation in (8) shows the corresponding representation via the meta-analytically pooled estimate $\hat{\theta}_{\text{MA}}$ and standard error $\sigma_{\text{MA}}$ [32]. While meta-analytic pooling could be extended to the random-effects model, this is typically not desired with only two studies for three reasons. First, the interest is in the true effects underlying the studies. Second, random-effects variance estimation is unreliable with only two studies. Finally, even if there is effect heterogeneity, fixed-effect meta-analysis is a valid procedure which estimates a well-defined average true effect [33].
@@ 3 Other p-value combination methods
While the two-trials rule and meta-analysis are the most commonly used p-value combination methods in practice, several other combination methods exist [17]. In this section, we examine Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods, which can also be used to obtain combined effect estimates, CIs, and p-values. Although these methods are not standard in drug regulation, they may have useful properties in certain settings, as we will demonstrate in the following.
@@ 3.1 Tippett’s (minimum) method
The combined p-value from Tippett’s method [34] is based on the minimum of the two p-values and given by
@@
It is closely related to the two-trials rule in the sense that the combined p-value based on the alternative “greater” from Tippett’s method is the same as one minus the combined p-value based on the alternative “less” from the two-trials rule, and vice versa [27]. Similarly, Tippett’s method is a special case of Wilkinson’s method based on the $r=1$ smallest out of $k=2$ p-values.
@@ 3.1.1 Effect estimation
The similarity to the two-trials rule is again visible as (10) looks similar to the estimation function from the two-trials rule (3) with the minimum and maximum flipped and using different normal quantiles. We can see that when the observed effect estimates are the same ($\hat{\theta}_{1}=\hat{\theta}_{2}=\hat{\theta}$), the median estimate ($a=1/2$) based on Tippett’s method is not equal to $\hat{\theta}$ but shifted from it, as the two-trials rule (see panels A and C in Figure 2 for an illustration). Similarly, CIs obtained from Tippett’s method are typically skewed in the sense that the distances between the point estimate and the upper and lower limits are not the same.
@@ 3.1.2 Asymptotics
that is, the more extreme of the two effects, see Appendix B.2 for details. In contrast to the two-trials rule, Tippett’s method is hence anti-conservative. This is illustrated in panel D of Figure 2 where Tippett’s CI is tightly concentrated around the larger effect estimate.
@@ 3.2 Fisher’s and Pearson’s (product) methods
Pearson’s and Fisher’s combination method are two closely related p-value combination methods, that are based on the product of p-values, or equivalently, the sum of the log p-values. Fisher’s method has been proposed for combining p-values from clinical trials [9, 35, 29], however, using the associated p-value function for effect estimation in a regulatory trials setting has remained unexplored.
The combined p-value function based on Fisher’s method [26] is given by
@@
while the combined p-value function based on Pearson’s method [36] is given by
@@
Pearson [37] proposed also another method based on the maximum of the test statistics underlying the p-values (11) and (12), but we will not consider this method here as its test statistic does not have an exact null distribution [38]. As with the two-trials rule and Tippett’s method, the combined p-value functions of Fisher’s and Pearson’s methods are related in the sense that the p-value function based on Fisher’s method and the alternative hypothesis “greater” is the same as one minus the p-value function based on Pearson’s method and the alternative “less”, and vice versa [27]. As we will show in the following, Fisher’s method also acts in a similar anti-conservative way as Tippett’s method, while Pearson’s method acts in a similar conservative way as the two-trials rule.
@@ 3.2.1 Effect estimation
$\hat{\mu}_{\text{P}}(a)=\begin{cases}\hat{\theta}+\sigma\,z_{\exp\{-\chi^{2}_{4}(a)/4\}}&\text{for alternative = "greater"}\\ \hat{\theta}-\sigma\,z_{\exp\{-\chi^{2}_{4}(a)/4\}}&\text{for alternative = "less"}\\ \end{cases}$ (14)
with $\chi^{2}_{4}(a)$ the $a\times 100\%$ quantile of the chi-squared distribution with four degrees of freedom. Importantly, the median estimates ($a=1/2$) from both methods do not equal the observed estimate $\hat{\theta}$ but are shifted away from it by $z_{\exp\{-\chi^{2}_{4}(1/2)/2\}}\approx-0.17$ standard errors $\sigma$, similar to the two-trials rule and Tippett’s method. Another similarity is that the CI is skewed since the distance between the lower and upper limits to the point estimate is not the same.
@@ 3.2.2 Asymptotics
see Appendix B.3 for details. This means that the combined estimation functions converge toward the more extreme effect for Fisher’s method (e.g., the maximum of two positively oriented effects), and the less extreme effect for Pearson’s method (e.g., the minimum of two positively oriented effects). The behavior is similar to Tippett’s method and the two-trials rule where one method acts anti-conservative (Fisher and Tippett’s methods), while the other methods acts conservative (Pearson’s method and the two-trials rule). However, the examples in panels B and D of Figure 2 suggest that in finite samples, Fisher’s and Pearson’s method remain closer to the weighted average compared to Tippett’s method and the two-trials rule.
@@ 3.3 Edgington’s (sum) method
Edgington’s method based on the sum of p-values [25] is yet another p-value combination method that can be used for obtaining a combined p-value function, and the last method that we will consider in this paper. It is given by
$p_{\text{E}}(\mu)=\begin{cases}E^{2}/2&\text{if}~{}0\leq E\leq 1\\ 1-(2-E)^{2}/2&\text{if}~{}1<E\leq 2\\ \end{cases}$ (15)
with $E=p_{1}(\mu)+p_{2}(\mu)$. An attractive feature is that two-sided CIs based on Edgington’s method are orientation invariant, which is not the case for the other combined p-value functions considered so far. That is, CIs based on Edgington’s method do not depend on the orientation of the underlying one-sided p-values, so the same CI is obtained regardless whether one uses p-values with the alternative ”greater” or ”less” [27]. Edgington’s method has previously been used in meta-analysis [27], to synthesize p-values from original and replication studies [39], and suggested as an alternative for the two-trials rule [22]. However, its estimation properties in the context of two trials remain unexplored.
@@ 3.3.1 Effect estimation
$\hat{\mu}_{\text{E}}(a)=\begin{cases}\hat{\theta}+\sigma\,z_{\sqrt{a/2}}&\text{for}~{}a\leq 1/2\\ \hat{\theta}-\sigma\,z_{\sqrt{(1-a)/2}}&\text{for}~{}a>1/2\\ \end{cases}$ (17)
for the alternative ”greater” and with the plus (minus) after $\hat{\theta}$ replaced with minus (plus) in (17) for the alternative ”less”. We can see that CIs obtained from (17) are symmetric and centered around the observed effect estimate $\hat{\theta}$, similar to meta-analysis but unlike the CIs from the two-trials rule, Fisher’s, and Pearson’s methods. Yet, Edgington’s CI is in this case narrower than the meta-analytic CI. For example, Edgington’s 95% CI is 12.2% narrower than the corresponding meta-analytic 95% CI. Panel A of Figure 2 illustrates this as Edgington’s CI is narrower than the meta-analytic CI, although both are centered around the same effect estimate. However, in case the trials’ effect estimates are different, Edgington’s CI can also be much wider. For instance, in panel B of Figure 2 where the trials produced very different results, Edgington’s CI is much wider than any of the other methods. This suggests that Edgington’s method reacts to heterogeneity by widening its CI to include both trial effect estimates.
@@ 3.3.2 Asymptotics
$\operatorname*{plim}_{\sigma_{1},\sigma_{2}\downarrow 0}\hat{\mu}_{\text{E}}(a)=\begin{cases}\min\{\theta_{1},\theta_{2}\}&\text{for}~{}a<1/2\\ \dfrac{\theta_{1}}{1+\sqrt{c}}+\dfrac{\theta_{2}}{1+1/\sqrt{c}}&\text{for}~{}a=1/2\\ \max\{\theta_{1},\theta_{2}\}&\text{for}~{}a>1/2\\ \end{cases}$ (18)
see Appendix B.3 for details. This means that CIs based on Edgington’s method will asymptotically always include both true effects, even when the trials’ sample sizes become arbitrarily large, see Figure 2D for an illustration. This behavior is strikingly different from meta-analysis whose CI shrinks to a point at the weighted average, even when the true effects are not the same.
@@ 4.1 The RESPIRE trials
We first revisit the RESPIRE trials [6, 7, 8], which were presented as motivating example in Table 1 in the introduction. The trials investigated the effect of ciprofloxacin in the treatment of non-cystic fibrosis bronchiectasis. Each trial had two treatment groups (on/off treatment cycles of either 14 or 28 days for 48 weeks) and two corresponding control groups. RESPIRE 1 showed a substantial treatment effect in the 14-day treatment regimen (estimated log rate ratio of $\log\widehat{\mathrm{RR}}=-0.49$ with 95% CI from $-0.85$ to 0.13)-0.13)- 0.13 ), while the benefit was less pronounced in RESPIRE 2 ($\log\widehat{\mathrm{RR}}=-0.18$ with 95% CI from $-0.53$ to $0.16$). Surprisingly, this was reversed for the 28-day regimens, with RESPIRE 2 showing a much stronger treatment effect ($\log\widehat{\mathrm{RR}}=-0.6$ with 95% CI from $-0.96$ to 0.23)-0.23)- 0.23 ) while RESPIRE 1 showed almost no benefit ($\log\widehat{\mathrm{RR}}=-0.02$ with 95% CI from $-0.39$ to 0.35)0.35)0.35 ). Figure 3 shows the p-value functions of the two studies (dashed lines) along with different combined p-value functions (solid lines) and corresponding point estimates and CIs (top). Table 2 shows the results in numerical form. Of note, all results were computed with our R package twotrials and Appendix A shows how the results for the 14-day treatment group can be reproduced.
Figure 3: Results of the RESPIRE trials [6, 7, 8] for the effect of ciprofloxacin over 14 days (top) or 28 days (bottom) compared to placebo for the treatment of non-cystic fibrosis bronchiectasis. The two-sided p-value functions of the individual trials (dashed lines), and the combined p-value functions (solid lines) based on the two-trials rule, fixed-effect meta-analysis, Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods are shown along with corresponding median estimates and CIs (95% and 99.875% via telescope lines). All p-values are based on the alternative ”greater” and then converted to two-sided p-values via the centrality function $2\min\{p,1-p\}$.
Table 2: Point estimates, 95% CIs, and p-values for the RESPIRE trials [6, 7, 8].
Lower CI Log rate ratio Upper CI CI width P-value (one-sided)
@@
   RESPIRE 1 $-0.85$ $-0.49$ $-0.13$ $0.72$ $0.00351$
     RESPIRE 2 $-0.53$ $-0.18$ $0.16$ $0.68$ $0.14400$
   Two-trials rule $-0.57$ $-0.28$ $-0.01$ $0.56$ $0.02073$
     Meta-analysis $-0.58$ $-0.33$ $-0.08$ $0.49$ $0.00432$
   Tippett $-0.68$ $-0.39$ $-0.08$ $0.59$ $0.00701$
     Fisher $-0.64$ $-0.35$ $-0.09$ $0.55$ $0.00434$
   Pearson $-0.58$ $-0.32$ $-0.04$ $0.53$ $0.01138$
     Edgington $-0.64$ $-0.34$ $-0.05$ $0.59$ $0.01088$
@@
   RESPIRE 1 $-0.39$ $-0.02$ $0.35$ $0.73$ $0.45699$
     RESPIRE 2 $-0.96$ $-0.60$ $-0.23$ $0.73$ $0.00064$
   Two-trials rule $-0.44$ $-0.12$ $0.17$ $0.61$ $0.20884$
     Meta-analysis $-0.57$ $-0.31$ $-0.05$ $0.52$ $0.00912$
   Tippett $-0.79$ $-0.50$ $-0.18$ $0.60$ $0.00127$
     Fisher $-0.75$ $-0.44$ $-0.12$ $0.62$ $0.00266$
   Pearson $-0.50$ $-0.18$ $0.13$ $0.62$ $0.12562$
     Edgington $-0.74$ $-0.31$ $0.12$ $0.86$ $0.10471$
@@
Looking at the CIs, we can see that meta-analysis produces narrower CIs than the other methods for both treatment regimens. The widest CIs are produced by Edgington’s method. For the 28-day regimen, Edgington’s CI is the only method that includes both trial effect estimates, and as a result is even wider than the CIs from the individual trials, reflecting the apparent heterogeneity. Looking at the decision based on the CIs, we can see that for the 14-day regimens, all 95% CIs exclude a log rate ratio of zero, the value of no effect, while all 99.875% CIs include it. However, for the 28-day regimens, the 95% CIs from meta-analysis, Fisher’s, and Tippett’s methods exclude zero. The other method’s 95% CIs include zero, but only Edgington’s method includes also the point estimate from RESPIRE 2. Finally, the 99.875% CIs of all methods include zero, thus leading to identical decisions at the one-sided $0.0006250.025^{2}=$ 0.000625 level. Note that for each method, the decision based on the CI is compatible with the combined p-values in Table 2, for example, a 99.875% CI excludes a log rate ratio of zero only if also the combined one-sided p-value is less than $0.000625$.

### v2

@@ 2.2 Fixed-effect meta-analysis (Stouffer’s method)
where
@@
and
$\sigma_{\text{MA}}=\frac{1}{\sqrt{1/\sigma_{1}^{2}+1/\sigma_{2}^{2}}}.$ (11)
The first equation in (9) represents Stouffer’s p-value combination method (after transforming p-values to $z$-values) using inverse standard errors as weights 19, whereas the second equation in (9) shows the corresponding representation via the meta-analytically pooled estimate (10) and standard error (11) 33. While meta-analytic pooling could be extended to the random-effects model, this is typically not desired with only two studies for three reasons. First, the interest is in the true effects underlying the studies. Second, random-effects variance estimation is unreliable with only two studies. Finally, even if there is effect heterogeneity, fixed-effect meta-analysis is a valid procedure which estimates a well-defined average true effect 34.
@@ 2.2.1 Effect estimation
The standard error (11) of the meta-analytic median estimate has two desirable properties: First, it is never larger than either of trials’ standard errors ($\sigma_{\text{MA}}\leq\min\{\sigma_{1},\sigma_{2}\}$). Second, under effect homogeneity, the standard error is the smallest among all unbiased estimators of the common effect 17. Both properties do not hold for the two-trials rule and the other p-value combination methods discussed below.
@@ 3 Other p-value combination methods
While the two-trials rule and meta-analysis are the most commonly used p-value combination methods in practice, several other combination methods exist 17. In this section, we examine Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods, which can also be used to obtain combined effect estimates, CIs, and p-values. Although these methods are not standard in drug regulation, they may have useful properties in certain settings, as we will demonstrate in the following.
@@ 3.1 Tippett’s (minimum) method
The combined p-value from Tippett’s method 35 is based on the minimum of the two p-values and given by
@@
It is closely related to the two-trials rule in the sense that the combined p-value based on the alternative “greater” from Tippett’s method is the same as one minus the combined p-value based on the alternative “less” from the two-trials rule, and vice versa 28. Similarly, Tippett’s method is a special case of Wilkinson’s method based on the $r=1$ smallest out of $k=2$ p-values.
@@ 3.1.1 Effect estimation
The similarity to the two-trials rule is again visible as (12) looks similar to the estimation function from the two-trials rule (3) with the minimum and maximum flipped and using different normal quantiles. In particular, the same median estimates are obtained (i.e., $\hat{\mu}_{\text{2TR}}(1/2)=\hat{\mu}_{\text{T}}(1/2)$) if opposite alternatives are specified.
We can see that when the observed effect estimates are the same ($\hat{\theta}_{1}=\hat{\theta}_{2}=\hat{\theta}$), the median estimate ($a=1/2$) based on Tippett’s method is not equal to $\hat{\theta}$ but shifted from it, as the two-trials rule (see panels A and C in Figure 2 for an illustration). Similarly, CIs obtained from Tippett’s method are typically skewed in the sense that the distances between the point estimate and the upper and lower limits are not the same.
@@ 3.1.2 Asymptotics
that is, the more extreme of the two effects, see Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation for details. In contrast to the two-trials rule, Tippett’s method is hence anti-conservative. This is illustrated in panel D of Figure 2 where Tippett’s CI is tightly concentrated around the larger effect estimate.
@@ 3.2 Fisher’s and Pearson’s (product) methods
Pearson’s and Fisher’s combination method are two closely related p-value combination methods, that are based on the product of p-values, or equivalently, the sum of the log p-values. Fisher’s method has been proposed for combining p-values from clinical trials 9, 36, 30, however, using the associated p-value function for effect estimation in a regulatory trials setting has remained unexplored.
The combined p-value function based on Fisher’s method 26 is given by
@@
while the combined p-value function based on Pearson’s method 37 is given by
@@
Pearson 38 proposed also another method based on the maximum of the test statistics underlying the p-values (13) and (14), but we will not consider this method here as its test statistic does not have an exact null distribution 39. As with the two-trials rule and Tippett’s method, the combined p-value functions of Fisher’s and Pearson’s methods are related in the sense that the p-value function based on Fisher’s method and the alternative hypothesis “greater” is the same as one minus the p-value function based on Pearson’s method and the alternative “less”, and vice versa 28. As we will show in the following, Fisher’s method also acts in a similar anti-conservative way as Tippett’s method, while Pearson’s method acts in a similar conservative way as the two-trials rule.
@@ 3.2.1 Effect estimation
$\hat{\mu}_{\text{P}}(a)=\begin{cases}\hat{\theta}-\sigma\,z_{\exp\{-\chi^{2}_{4}(a)/4\}}&\text{for alternative = "greater"}\\ \hat{\theta}+\sigma\,z_{\exp\{-\chi^{2}_{4}(a)/4\}}&\text{for alternative = "less"}\\ \end{cases}$ (16)
with $\chi^{2}_{4}(a)$ the $a\times 100\%$ quantile of the chi-squared distribution with four degrees of freedom. Importantly, the median estimates ($a=1/2$) from both methods do not equal the observed estimate $\hat{\theta}$ but are shifted away from it by $z_{\exp\{-\chi^{2}_{4}(1/2)/4\}}\approx-0.17$ standard errors $\sigma$, similar to the two-trials rule and Tippett’s method. Another similarity is that the CI is skewed since the distance between the lower and upper limits to the point estimate is not the same.
@@ 3.2.2 Asymptotics
see Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation for details. This means that the combined estimation functions converge toward the more extreme effect for Fisher’s method (e.g., the maximum of two positively oriented effects), and the less extreme effect for Pearson’s method (e.g., the minimum of two positively oriented effects). The behavior is similar to Tippett’s method and the two-trials rule where one method acts anti-conservative (Fisher and Tippett’s methods), while the other methods acts conservative (Pearson’s method and the two-trials rule). However, the examples in panels B and D of Figure 2 suggest that in finite samples, Fisher’s and Pearson’s method remain closer to the weighted average compared to Tippett’s method and the two-trials rule.
@@ 3.3 Edgington’s (sum) method
Edgington’s method based on the sum of p-values 25, 40 is yet another p-value combination method that can be used for obtaining a combined p-value function, and the last method that we will consider in this paper. It is given by
$p_{\text{E}}(\mu)=\begin{cases}E^{2}/2&\text{if}\leavevmode\nobreak\ 0\leq E\leq 1\\ 1-(2-E)^{2}/2&\text{if}\leavevmode\nobreak\ 1<E\leq 2\\ \end{cases}$ (17)
with $E=p_{1}(\mu)+p_{2}(\mu)$. An attractive feature is that two-sided CIs based on Edgington’s method are orientation invariant, which is not the case for the other combined p-value functions considered so far. That is, CIs based on Edgington’s method do not depend on the orientation of the underlying one-sided p-values, so the same CI is obtained regardless whether one uses p-values with the alternative "greater" or "less" 28. Edgington’s method has previously been used in meta-analysis 28, to synthesize p-values from original and replication studies 40, and suggested as an alternative for the two-trials rule 22. However, its estimation properties in the context of two trials remain unexplored.
@@ 3.3.1 Effect estimation
The standard error of Edgington’s median estimate is given by
$\sigma_{\text{E}}=\frac{\sqrt{2}}{1/\sigma_{1}+1/\sigma_{2}}$ (19)
and does not depend on the true effects, similar to meta-analysis but unlike the two-trials rule. It is always larger than the meta-analytic standard error (11), see Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation. Therefore, under effect homogeneity, Edgington’s method is less efficient than meta-analysis at estimating the common effect. Under effect heterogeneity, however, the two methods target different estimands, so a comparison of their standard errors may not be meaningful. Finally, unlike meta-analysis, Edgington’s standard error is not always smaller or equal to either of the two trials’ standard errors. This is only the case if the standard error ratio is $\sqrt{2}-1\leq\sigma_{2}/\sigma_{1}\leq\sqrt{2}+1$. For example, suppose $\sigma_{1}=0.5$ and $\sigma_{2}=2$, then Edgington’s standard error is $\sqrt{2}/(2+0.5)=0.566$, which is greater than $\sigma_{1}$
@@
$\hat{\mu}_{\text{E}}(a)=\begin{cases}\hat{\theta}+\sigma\,z_{\sqrt{a/2}}&\text{for}\leavevmode\nobreak\ a\leq 1/2\\ \hat{\theta}-\sigma\,z_{\sqrt{(1-a)/2}}&\text{for}\leavevmode\nobreak\ a>1/2\\ \end{cases}$ (20)
for the alternative "greater" and with the plus (minus) after $\hat{\theta}$ replaced with minus (plus) in (20) for the alternative "less". We can see that CIs obtained from (20) are symmetric and centered around the observed effect estimate $\hat{\theta}$, similar to meta-analysis but unlike the CIs from the two-trials rule, Tippett’s, Fisher’s, and Pearson’s methods. Yet, Edgington’s CI is in this case narrower than the meta-analytic CI. For example, Edgington’s 95% CI is 12.2% narrower than the corresponding meta-analytic 95% CI. Panel A of Figure 2 illustrates this as Edgington’s CI is narrower than the meta-analytic CI, although both are centered around the same effect estimate. However, in case the trials’ effect estimates are different, Edgington’s CI can also be much wider. For instance, in panel B of Figure 2 where the trials produced very different results, Edgington’s CI is much wider than any of the other methods. This suggests that Edgington’s method reacts to heterogeneity by widening its CI to include both trial effect estimates.
@@ 3.3.2 Asymptotics
$\operatorname*{plim}_{\sigma_{1},\sigma_{2}\downarrow 0}\hat{\mu}_{\text{E}}(a)=\begin{cases}\min\{\theta_{1},\theta_{2}\}&\text{for}\leavevmode\nobreak\ a<1/2\\ \dfrac{\theta_{1}}{1+\sqrt{c}}+\dfrac{\theta_{2}}{1+1/\sqrt{c}}&\text{for}\leavevmode\nobreak\ a=1/2\\ \max\{\theta_{1},\theta_{2}\}&\text{for}\leavevmode\nobreak\ a>1/2\\ \end{cases}$ (21)
see Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation for details. This means that CIs based on Edgington’s method will asymptotically always include both true effects, even when the trials’ sample sizes become arbitrarily large, see panel D of Figure 2 for an illustration. This behavior is strikingly different from meta-analysis whose CI shrinks to a point at the weighted average, even when the true effects are not the same.
@@ 4.1 The RESPIRE trials
We first revisit the RESPIRE trials 6, 7, 8, which were presented as motivating example in Table 1 in the introduction. The trials investigated the effect of ciprofloxacin in the treatment of non-cystic fibrosis bronchiectasis. Each trial had two treatment groups (on/off treatment cycles of either 14 or 28 days for 48 weeks) and two corresponding control groups. RESPIRE 1 showed a substantial treatment effect in the 14-day treatment regimen (estimated log rate ratio of $\log\widehat{\mathrm{RR}}=-0.49$ with 95% CI from $-0.85$ to 0.13)-0.13)- 0.13 ), while the benefit was less pronounced in RESPIRE 2 ($\log\widehat{\mathrm{RR}}=-0.18$ with 95% CI from $-0.53$ to $0.16$). Surprisingly, this was reversed for the 28-day regimens, with RESPIRE 2 showing a much stronger treatment effect ($\log\widehat{\mathrm{RR}}=-0.6$ with 95% CI from $-0.96$ to 0.23)-0.23)- 0.23 ) while RESPIRE 1 showed almost no benefit ($\log\widehat{\mathrm{RR}}=-0.02$ with 95% CI from $-0.39$ to 0.35)0.35)0.35 ). Figure 3 shows the p-value functions of the two studies (dashed lines) along with different combined p-value functions (solid lines) and corresponding point estimates and CIs (top). Table 2 shows the results in numerical form. Of note, all results were computed with our R package twotrials and Appendix Combined P-value Functions for Compatible Effect Estimation and Hypothesis Testing in Drug Regulation shows how the results for the 14-day treatment group can be reproduced.
Figure 3: Results of the RESPIRE trials 6, 7, 8 for the effect of ciprofloxacin over 14 days (top) or 28 days (bottom) compared to placebo for the treatment of non-cystic fibrosis bronchiectasis. The two-sided p-value functions of the individual trials (dashed lines), and the combined p-value functions (solid lines) based on the two-trials rule, fixed-effect meta-analysis, Tippett’s, Fisher’s, Pearson’s, and Edgington’s methods are shown along with corresponding median estimates and CIs (95% and 99.875% via telescope lines). All p-values are based on the alternative "greater" and then converted to two-sided p-values via the centrality function $2\min\{p,1-p\}$.
Table 2: Point estimates (with implicit weights), 95% CIs (with widths), and p-values for the RESPIRE trials 6, 7, 8. Note that weigths and CI widths are computed from unrounded numbers and may not always exactly correspond to the rounded point estimates and CIs.
Log rate ratio Weight RESPIRE 1 95% CI CI width P-value (one-sided)
@@
[…]
@@
[…]
@@
It is interesting to consider the median estimate as a weighted average of the trial specific point estimates, and to determine the corresponding (implicit) weights. Table 2 reports the weight of the point estimate from RESPIRE 1 toward the median estimate (the weight from RESPIRE 2 is one minus the weight from RESPIRE 1). The more extreme estimate (RESPIRE 1 in the 14-day group, and RESPIRE 2 in the 28-day group) contributes more to Tippett’s and Fisher’s estimates and less to Pearson’s and the two-trials rule estimates, which aligns with the expected behavior. Similarly, the weight of RESPIRE 1 is slightly larger for Edgington’s method than meta-analysis because Edgington’s estimate gives more weights to trials with larger standard errors due to its inverse standard error weighting.
Looking at the CIs, we can see that meta-analysis produces narrower CIs than the other methods for both treatment regimens. The widest CIs are produced by Edgington’s method. For the 28-day regimen, Edgington’s 95% CI is the only method that includes both trial effect estimates, and as a result is even wider than the CIs from the individual trials, reflecting the apparent heterogeneity. Looking at the decision based on the CIs, we can see that for the 14-day regimens, all 95% CIs exclude a log rate ratio of zero, the value of no effect, while all 99.875% CIs include it. However, for the 28-day regimens, the 95% CIs from meta-analysis, Fisher’s, and Tippett’s methods exclude zero. The other method’s 95% CIs include zero, but only Edgington’s method includes also the point estimate from RESPIRE 2. Finally, the 99.875% CIs of all methods include zero, thus leading to identical decisions at the one-sided $0.0006250.025^{2}=$ 0.000625 level. Note that for each method, the decision based on the CI is compatible with the combined p-values in Table 2, for example, a 99.875% CI excludes a log rate ratio of zero only if also the combined one-sided p-value is less than $0.000625$.